OpenStudy (hannahp):

Heeelllpp PLEASE! 5(2n + 1) + -5 = 10n True or False? (7p) 1/7 + - p = 0 True or False?

OpenStudy (turingtest):

Want how to do it or just answers?

OpenStudy (anonymous):

n Π ... [ 1/(2i+1)(2i+2) ] i=0 = [1/(1·2) ] • [1/(3·4)] • [1/(5·6)] • ··· • [ 1 / (2n+1)(2n+2) ] = 1 / [ 1·2·3·4·5·6· ... · (2n+1)· (2n+2) ] = 1 / (2n+2)! _____________________________________ Hence, the result in question is TRUE. ... Ans. ______________________________________

OpenStudy (hannahp):

The Second One Is: